OpenStudy (anonymous):

Perform the operations. Write all answers in the form of a+bi. 1. (2-7i)(-3+4i) 2. (7-sqrt-9)-(4+sqrt-4)

OpenStudy (anonymous):

2. \[(7-\sqrt{-9})-(4+\sqrt{-4})\]

OpenStudy (anonymous):

Have you heard of FOIL??

OpenStudy (anonymous):

Yes. I've already done it i was just hoping someone else could go through it to make sure I did it correctly.

OpenStudy (anonymous):

It will be more appreciable if you show your work to us... So that we can check..

OpenStudy (anonymous):

Typically I learn better if I can see the steps that someone else went through, that way if it is wrong at least I learned how to do it. 1. I got (2-7i)(-3+4i)= -6+2*4i+3*7i-28i= -6+8i+21i-28i= -6+i= -6+0i #2 seems to have run away from me so I'll have to find that one.

OpenStudy (anonymous):

Can you tell: \[\large (-7i)(+4i) = ??\]

OpenStudy (anonymous):

And typically if you have faith on what you did then you will not look at others work what they did in solving the same..

OpenStudy (anonymous):

I don't have faith in what I did, at all. That's why I'm here. But that doesn't change how I learn.

OpenStudy (anonymous):

Right or wrong?? This does not matter.. What matter is Having Faith on ourselves.. Of course you are here for learning but you can't learn without having faith on yourself.. And that is for sure... Now can you please tell me: \[\large (-7i)(+4i) = ??\]

OpenStudy (anonymous):

(2-7i)(-3+4i) -6+8i+21i-28i^2 29i-6-28x-1 29i-6+28 29i+22

OpenStudy (anonymous):

I^2=-1

OpenStudy (anonymous):

(7-sqrt(-9))-(4+sqrt(-4)) (7-3i)-(4+2i) 3-5i

OpenStudy (anonymous):

I think @StefaniTriscuit is more busy helping others.. Ha ha ha.. Ignoring his or her own question..